OpenStudy (anonymous):

Find the product (2n + 2)(2n - 2)

OpenStudy (ajprincess):

\((a+b)(a-b)=a^2-b^2\) Here a=2n, b=2 Does that help?

OpenStudy (anonymous):

is it 4n^2 - 4?

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

so if the question is find the square (6m + 7)^2 what formula is it?

OpenStudy (ajprincess):

\((a+b)^2=a^2+2ab+b^2\) here a=6m, b=7

OpenStudy (ajprincess):

Does that help?

OpenStudy (anonymous):

woah i think you just taught me math....

OpenStudy (anonymous):

so its 36m^3 + 84m + 49?

OpenStudy (ajprincess):

it is 36m^2+ 84m + 49

OpenStudy (anonymous):

well 36m^2 i mean

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

so i was right?

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

this is the first time I have done algebra ever really, so thanks for the help!

OpenStudy (anonymous):

if its (2x - 6)^2 what formula?

OpenStudy (anonymous):

"find the square"

OpenStudy (ajprincess):

\((a-b)^2=a^2-2ab+b^2\)

OpenStudy (ajprincess):

here a=2x, b=6

OpenStudy (anonymous):

working on it 1sec

OpenStudy (anonymous):

is it 4x^2 - 8x + 36?

OpenStudy (ajprincess):

middle term is wrong

OpenStudy (anonymous):

hmm its either 12x or 24x not sure

OpenStudy (anonymous):

id say 24x?

OpenStudy (ajprincess):

2*2x*6=?

OpenStudy (anonymous):

(2n)^2-(2)^2 4n^2-4

OpenStudy (anonymous):

so it is 24x!

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

Do you know how to find missing coefficient?

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

ok (4d - 4)(3d - 2) = 12d^2 + ?d + 8 ? = missing coefficient

OpenStudy (ajprincess):

(4d - 4)(3d - 2) =4d(3d-2)-4(3d-2) =12d^2-8d-12d+8

OpenStudy (ajprincess):

?d=-8d-12d

OpenStudy (anonymous):

ohhhh i see so 4

OpenStudy (ajprincess):

nope:(

OpenStudy (ajprincess):

both the numbers are negative. so the result is also negative

OpenStudy (anonymous):

blah its okay, How would i factor the expression 3x^3 + 3x^2 + x + 1

OpenStudy (anonymous):

-20?

OpenStudy (anonymous):

Yeah I should have known better